OpenStudy (anonymous):

Arc BC and arc DE both measure 70°. Arc CD measures 160°. What is the measure of angle BAE? A. 60 degrees B. 120 degrees C. 50 degrees D. 120 degrees

OpenStudy (amistre64):

i think there might need to be a picture attached to this; or a crudly drawn facsimile

OpenStudy (anonymous):

Yeah, I wasn't exactly sure if there should be. I'll try and attach the file or draw it haha, thanks :)

OpenStudy (amistre64):

 im thinking itll look like this; but i cant be sure

OpenStudy (anonymous):

yeah that's how it looks :)

OpenStudy (amistre64):

theres a thrm that says; the (outerarc - inner arc) divided in half give us the angle we want

OpenStudy (amistre64):



OpenStudy (amistre64):

you are given the outer, and the sides; so determine the inner from what you know; the missing inner is the difference from 360 degrees

OpenStudy (anonymous):

So would I have to add 160 and 70? and then subtract it from 360?

OpenStudy (anonymous):

or I mean 160 and 140?

OpenStudy (amistre64):

160+140 = 300 360 - 300 = 60 ; your inner arc = 60 160 - 60 = 100 ; 100/2 = 50 for the angle

OpenStudy (amistre64):

so yes :)

OpenStudy (anonymous):

haha, okay! thank you so much :)

OpenStudy (amistre64):

yw, sorry about the delay ut internet went down at the school

OpenStudy (anonymous):

haha no problem :)